# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ Abstract
The control of interactions among quantum emitters through nanophotonic structures offers significant potential for quantum technologies. However, a theoretical description of the interaction of multiple quantum emitters with complex dispersive dielectric objects remains highly challenging. Here we introduce an approach based on the modified Langevin noise formalism that unveils the roles of both the noise polarization currents of the dielectrics and the vacuum fluctuations of the electromagnetic field scattered by the dielectrics. This extends Refs. [1], [2] to the general case of an arbitrary number of emitters. The proposed approach allows us to describe the dynamics of the quantum emitters for arbitrary initial quantum states of the electromagnetic environment consisting of two independent bosonic reservoirs, a medium-assisted reservoir and a scattering-assisted reservoir, each characterized by its own spectral density matrix. Understanding how these reservoirs shape emitter dynamics is crucial to understanding light–matter interactions in complex electromagnetic environments and to enhancing intrinsic emitter properties in structured environments.
@@ I Introduction
The design of qubit-qubit interactions is fundamental to a wide range of quantum technologies, including quantum networking, quantum information processing, and quantum computation (e.g., [3], [4], [5], [6]). Nanophotonics enables the engineering and control of quantum properties of light by embedding quantum emitters within nanostructured environments. In particular, systems that utilize quantum emitters as qubits have attracted growing interest [7, 8], with significant efforts devoted to tailoring their mutual interactions through metallic or dielectric nano-structures [9]. These systems are physically rich: metal and dielectric nanoparticles, typically dispersive and lossy, can support plasmonic and dielectric resonances of various multipolar orders, which can be engineered by adjusting the shape and spatial arrangement of the nanoparticles, or by exploiting collective resonances. Moreover, nanophotonic devices enable subwavelength confinement of light, allowing device miniaturization, enhanced light–matter interactions, and consequently faster dynamics and higher operating speeds. Such engineered couplings play a central role in the generation and manipulation of non-classical states of light. Nevertheless, providing a theoretical description of multiple quantum emitters that interact with realistic nanostructures remains highly nontrivial. The electromagnetic environment is inherently dispersive, and its modes span a high-dimensional continuum. The account of the finite extent, material losses, and dispersion of the dielectric components poses serious challenges to the canonical quantization of the electromagnetic field. In this context, macroscopic quantum electrodynamics (e.g. [10], [11]), offers a widely adopted phenomenological framework, enabling a consistent and flexible quantization scheme in the presence of complex dielectric media.
@@
Macroscopic quantum electrodynamics relies on the Langevin noise formalism, which is based on the fluctuation dissipation theorem. The electromagnetic field, commonly referred to as the medium-assisted field, emerges from the dielectric noise polarization current, mediated by the dyadic Green function of the dielectric objects. As argued in refs. [12] and [13] the original macroscopic quantum electrodynamic model disregards the influence of vacuum fluctuations of the electromagnetic field scattered by the dielectric objects, called the scattering-assisted field in [14] and [15]. The modified Langevin noise formalism adds the scattering-assisted field to the medium-assisted field: polarization current fluctuations and vacuum fluctuations of the electromagnetic field scattered by the dielectric objects are on the same footing. Recently, it has been justified for finite-size dispersive dielectric objects with arbitrary shapes applying the Heisenberg picture to a phenomenological model of dielectric media based on a continuum set of harmonic oscillators [15]. Each elementary region of the dielectric medium is described as a bath of harmonic oscillators that couples to the electromagnetic field in such a way that the quantized electromagnetic field experiences the dielectric medium through its macroscopic dielectric permittivity (e.g., [16], [17]). As the medium-assisted field, the scattering-assisted field can also be expressed in terms of vacuum fluctuations through the dyadic Green function of the dielectric objects [15].
@@
The modified Langevin noise formalism has recently been applied for the first time to a single quantum emitter that interacts with a linear dispersive dielectric slab in [14], [1], and with a linear dispersive dielectric sphere in [2]. The quantum emitter interacts with two distinct and independent bosonic reservoirs, the medium-assisted reservoir and the scattering-assisted reservoir, each characterized by its own spectral density. The medium-assisted and the scattering-assisted spectral densities are quadratic functionals of the dyadic Green function of dielectric objects $\mathcal{G}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})$. The time evolution of the reduced density operator of the quantum emitter depends on both the initial quantum state of the entire system and the spectral densities of the two reservoirs. Only under specific conditions can the actions of the two reservoirs combine in such a way that they can be represented by a single equivalent bosonic reservoir characterized by an equivalent spectral density depending only on the value of the dyadic Green function of the dielectric object at the quantum emitter position. In fact, when the initial quantum state of the entire system is a product state and both reservoirs are initially in the vacuum state, the reduced-density operator of the quantum emitter evolves as the emitter interacts with an equivalent single bosonic reservoir, with only positive frequencies and equivalent spectral density $\mathcal{J}(\omega)$ equal to the sum of the spectral density of the individual reservoirs. In this case, it results that [14], [1], [2], $\mathcal{J}(\omega)=\frac{\omega^{2}}{\pi\hbar\varepsilon_{0}c^{2}}\mathbb{p}\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{a},\mathbf{r}_{a}\right)\right]\cdot\mathbb{p}$ as a consequence of the fundamental integral relation 31 of ref. [15]; here $\mathbf{r}_{a}$ is the position vector of the emitter and $\mathbb{p}$ is its transition dipole moment. When the two reservoirs are initially in thermal quantum states with equal temperature $T_{0}$ the reduced dynamics of the quantum emitter can yet be described by a single zero-temperature bosonic reservoir with positive and negative frequencies and an equivalent temperature-dependent spectral density equal to $\theta(\omega,T_{0})\mathcal{J}(|\omega|)$ where $\theta(\omega,T_{0})=\frac{1}{2}sign(\omega)\left[1+\coth\left(\frac{\beta_{0}\hbar\omega}{2}\right)\right]$ and $\beta_{0}=1/(k_{B}T_{0})$. In the literature based on the Langevin noise formalism, similar relations are widely used; however, the conditions under which they remain valid are not always clearly stated. Indeed, if the two reservoirs have different temperatures, simply knowing the values of the dyadic Green function of the dielectric objects at the position of the quantum emitter is not sufficient to evaluate the dynamics of the emitter. In this case, the reduced dynamics of the emitter can yet be described by a zero temperature single bosonic reservoir with an equivalent spectral density equal to the sum of the temperature-dependent medium-assisted spectral density and the temperature-dependent scattering-assisted spectral density [2]. In general, if the initial states of the two reservoirs are not thermal states, the situation becomes considerably more complicated, and it is no longer possible to describe their action through an equivalent single reservoir.
@@
Although the model introduced in refs. [1], [2] is capable of handling an electromagnetic environment with arbitrary initial quantum state, the considered framework is limited to a single quantum emitter. In the present article, we lift this restriction and generalize the approach based on the modified Langevin noise formalism to a collection of multiple quantum emitters with arbitrary orientations of the transition dipole moments. As we did for the single quantum emitter [1], we applied the emitter-centered mode technique to reduce the number of degrees of freedom of the electromagnetic environment (e.g. [18], [19]). In particular, we show also in this case that the reduced dynamic of the quantum emitters depends only on the values of the dyadic Green function at the quantum emitter positions if the initial state of the entire system is a product state and the initial states of the medium-assisted and the scattering-assisted reservoirs are thermal states with the same temperature. When these conditions are not satisfied, the reduced dynamics of the quantum emitter depends on the generalized spectral densities of the single reservoirs, which are quadratic functionals of the dyadic Green function of the dielectric objects. We first introduce the generalized spectral density matrices that characterize the medium-assisted and the scattering-assisted reservoirs and discuss their general properties. Then, we introduce an equivalent reduced electromagnetic environment characterized by a single spectral density matrix for scenarios in which the two bosonic reservoirs are initially in thermal quantum states with different temperatures. We eventually applied the approach to study the correlation between two quantum emitters interacting with a Drude sphere or with a Drude nanostructure composed of two rods and a disk. We first compute the medium-assisted and the scattering-assisted spectral density matrices for the two cases, then we evaluate the evolution of the reduced density operator of the two quantum emitters and find entanglement decay and revivals, as well as entanglement generation.
@@
In Section II we apply the modified Langevin noise formalism to model multiple quantum emitters interacting with a dispersive electromagnetic environment. In Section III, we use emitter-centered modes to reduce the number of degrees of freedom of the electromagnetic environment. In Section IV we introduce the medium- and the scattering-assisted spectral density matrices. In Section V, we propose a surrogate bosonic environment to describe the reduced dynamics of quantum emitters when the initial quantum state of the entire system is a product state, and the medium-assisted and scattering-assisted bosonic reservoirs are initially in thermal quantum states with different temperatures. In Section VI we apply our model to a system of two quantum emitters. In Section VII, we give a summary and discuss our conclusions.
@@ II Model
A collection of $N$ quantum emitters mutually interacts and couples to finite-size dielectric objects of arbitrary shape embedded in an unbounded space. We denote by $\mathbf{r}_{j}$ the position vector of the $j$-th quantum emitter with j=1,2,,Nj=1,2,...,N. We assume that the dielectrics are linear, isotropic, and dispersive in time. We denote by $V$ the region occupied by the dielectric objects and by $\varepsilon_{\omega}(\mathbf{r})$ their relative dielectric permittivity in the frequency domain. The dispersive dielectrics together with the electromagnetic field constitute the electromagnetic environment of the quantum emitters.
@@ II.1 Hamiltonian
The Hamiltonian of the entire system, multiple quantum emitters ++ electromagnetic environment, is given by
@@
where $\hat{H}_{A}$ is the bare Hamiltonian of the quantum emitters, assumed to be mutually isolated, $\hat{H}_{E}$ is the bare Hamiltonian of the electromagnetic environment, and $\hat{H}_{I}$ describes their interaction.
In the multipolar coupling scheme (Power–Zienau–Woolley picture) and within the dipole approximation the interaction Hamiltonian $\hat{H}_{I}$ is given by
$\hat{H}_{I}=-\sum_{j=1}^{N}\hat{\mathbb{d}}_{j}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j})$ (2)
where $\hat{\mathbf{E}}(\mathbf{r})$ is the electric field operator and $\hat{\mathbb{d}}_{j}$ is the electric dipole moment operator of the $j$-th emitter located at $\mathbf{r}_{j}$. We assume that the dipole moment operator of the $j$-th emitter couples through a fixed polarization direction $\mathbb{u}_{j}$ as in ref. [19], so that $\hat{\mathbb{d}}_{j}=\hat{\mu}_{j}\mathbb{u}_{j}$ where $\hat{\mu}_{j}$ is the corresponding transition dipole moment operator. More generally, one may retain up to three orthogonal polarization components per emitter, $\hat{\mathbb{d}}_{j}=\sum_{a}\hat{\mu}_{ja}\mathbb{a}$ where $a=x,y,z$, and $\hat{\mu}_{jx},\hat{\mu}_{jy},\hat{\mu}_{jz}$ are the corresponding transition-dipole moment operators.
II.2 Diagonal Form of the Electromagnetic Environment Hamiltonian
The bare Hamiltonian of the electromagnetic environment $\hat{H}_{E}$ accounts for the electromagnetic field, the polarization currents of the dielectric objects, and their interaction. The modified Langevin noise formalism provides a straightforward way to diagonalize it [15].
We express the electric field operator in the Schrödinger picture as $\hat{\mathbf{E}}(\mathbf{r})=\int^{\infty}_{0}d\omega[\hat{\mathbf{E}}_{\omega}(\mathbf{r})+h.c.]$ where $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ is its monochromatic component in the Heisenberg picture. In the modified Langevin noise formalism, the electric field operator $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ has two contributions, the medium-assisted contribution $\hat{\mathbf{E}}^{(M)}_{\omega}(\mathbf{r})$ and the scattering-assisted contribution $\hat{\mathbf{E}}^{(S)}_{\omega}(\mathbf{r})$,
@@ II.2 Diagonal Form of the Electromagnetic Environment Hamiltonian
Let $\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})$ denote the solution of homogeneous equation
$\left(\nabla_{\mathbf{r}}\times\nabla_{\mathbf{r}}\times-k_{\omega}^{2}\varepsilon_{\omega}\right)\mathbf{F}_{\omega\mathbf{n}\nu}=0$ (7)
obeying the boundary condition
$\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})\underset{r\rightarrow\infty}{\approx}e^{ik_{\omega}\mathbf{r}\cdot\mathbf{n}}\mathbb{e}_{\mathbf{n}\nu},$ (8)
@@
where $\hat{g}_{\omega\mathbf{n}\nu}$ is the monochromatic bosonic operator that describes the fluctuation of the radiation incoming from infinity and scattered by the dielectric object. Here $o_{\mathbf{n}}=(\theta_{\mathbf{n}},\phi_{\mathbf{n}})$ are the polar angles of the unit vector $\mathbf{n}$}, $do_{\mathbf{n}}=\sin\theta_{\mathbf{n}}d\theta_{\mathbf{n}}d\phi_{\mathbf{n}}$ is the differential of the solid angle and the integration is performed over the whole solid angle with $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$.
@@
${\left[\hat{\mathbf{f}}_{\omega}(\mathbf{r}),\hat{\mathbf{f}}_{\omega^{\prime}}^{\dagger}\left(\mathbf{r}^{\prime}\right)\right]}$ $=\delta\left(\omega-\omega^{\prime}\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)I,$ (11a)
@@
where $\delta\left(o_{\mathbf{n}}-o_{\mathbf{n}^{\prime}}\right)=\delta\left(\theta_{\mathbf{n}}-\theta_{\mathbf{n}}^{\prime}\right)\delta\left(\varphi_{\mathbf{n}}-\varphi_{\mathbf{n}}^{\prime}\right)/\sin\theta_{\mathbf{n}}$. These commutation relations, the expression of the medium-assisted electric field $\hat{\mathbf{E}}^{(M)}_{\omega}$ and the expression of the scattering-assisted electric field $\hat{\mathbf{E}}^{(S)}_{\omega}$ guarantee the canonical commutation relations for the electromagnetic field and the bath oscillator fields describing the dielectric objects [15].
@@
The fundamental integral identity [15]

### v2

@@ Abstract
The control of interactions among quantum emitters through nanophotonic structures offers significant opportunities for quantum technologies. However, a rigorous theoretical description of the interaction of multiple quantum emitters with complex, dispersive dielectric objects remains challenging. Here, we introduce an approach based on the modified Langevin noise formalism that unveils the roles of both the noise polarization currents of the dielectrics and the vacuum fluctuations of the electromagnetic field scattered by the dielectrics. This work extends Refs. [1] and [2] to the general case of an arbitrary number of emitters. The proposed approach allows us to describe the dynamics of the quantum emitters for arbitrary initial quantum states of the electromagnetic environment, consisting of two independent bosonic reservoirs, a medium-assisted reservoir and a scattering-assisted reservoir, each characterized by its own spectral density matrix. Specifically, we examine situations where both reservoirs are initially in thermal quantum states but have different temperatures. Understanding how these reservoirs shape the dynamics of the emitters is crucial for understanding light–matter interactions in complex electromagnetic environments and for improving intrinsic emitter properties within structured environments.
@@ I Introduction
The design of qubit-qubit interactions is fundamental to a wide range of quantum technologies, including quantum networking, quantum information processing, and quantum computation (e.g., [3], [4], [5], [6]). Nanophotonics enables the engineering and control of quantum properties of light by embedding quantum emitters within nanostructured environments. In particular, systems that utilize quantum emitters as qubits have attracted growing interest [7, 8], with significant efforts devoted to tailoring their mutual interactions through metallic or dielectric nanostructures [9]. These systems are physically rich: metal and dielectric nanoparticles, typically dispersive and lossy, can support plasmonic and dielectric resonances of various multipolar orders, which can be engineered by adjusting the shape and spatial arrangement of the nanoparticles, or by exploiting collective resonances. Moreover, nanophotonic devices enable subwavelength confinement of light, allowing device miniaturization, enhanced light–matter interactions, and consequently faster dynamics and higher operating speeds. Such engineered couplings play a central role in the generation and manipulation of non-classical states of light. Nevertheless, providing a theoretical description of multiple quantum emitters that interact with realistic nanostructures remains highly nontrivial. The electromagnetic environment is inherently dispersive, and its modes span a high-dimensional continuum. Accounting for the finite extent, material losses, and dispersion of the dielectric components poses serious challenges to the canonical quantization of the electromagnetic field. In this context, macroscopic quantum electrodynamics (e.g. [10], [11]), offers a widely adopted phenomenological framework, enabling a consistent and flexible quantization scheme in the presence of complex dielectric media.
@@
Macroscopic quantum electrodynamics relies on the Langevin noise formalism, which is based on the fluctuation-dissipation theorem. The electromagnetic field, commonly referred to as the medium-assisted field, emerges from the dielectric noise polarization current, mediated by the dyadic Green function of the dielectric objects. As argued in [12] and [13], the original macroscopic quantum electrodynamic model disregards the influence of vacuum fluctuations of the electromagnetic field scattered by the dielectric objects, called the scattering-assisted field in [14] and [15]. The modified Langevin noise formalism adds the scattering-assisted field to the medium-assisted field: polarization current fluctuations and vacuum fluctuations of the electromagnetic field scattered by the dielectric objects are on the same footing. This framework was justified on a rigorous theoretical basis by Ciattoni [15], who derived it for finite, dispersive dielectric bodies of arbitrary shape by formulating the theory in the Heisenberg picture within a phenomenological model of dielectric media based on a continuum of harmonic oscillators. Each elementary region of the dielectric medium is described as a continuum of harmonic oscillators that couples to the electromagnetic field, so that the quantized field experiences the dielectric medium through its macroscopic dielectric permittivity (e.g., [16], [17]). As the medium-assisted field, the scattering-assisted field can also be expressed in terms of vacuum fluctuations through the dyadic Green function of dielectric objects [15].
@@
The modified Langevin noise formalism has recently been applied, for the first time, to a single quantum emitter interacting with a linear, dispersive dielectric slab in [14], [1], and with a linear, dispersive dielectric sphere in [2]. In this framework, the quantum emitter couples to two distinct and independent bosonic reservoirs: the medium-assisted reservoir and the scattering-assisted reservoir. Each reservoir is characterized by its own spectral density, and both spectral densities are quadratic functionals of the dyadic Green function of the dielectric objects, $\mathcal{G}_{\omega}(\mathbf{r},\mathbf{r}^{\prime})$. The time evolution of the reduced density operator of the emitter depends on both the initial quantum state of the entire system and on the spectral densities of the two reservoirs. Only under specific conditions can the actions of the two reservoirs be represented by a single equivalent bosonic reservoir whose spectral density depends solely on the value of the dyadic Green function of the dielectric object at the quantum emitter position. In fact, when the initial quantum state of the entire system is a product state and both reservoirs are initially in the vacuum state, the reduced density operator of the quantum emitter evolves as the emitter interacts with an equivalent single bosonic reservoir, with only positive frequencies and equivalent spectral density $\mathcal{J}(\omega)$ equal to the sum of the spectral density of the individual reservoirs. In this case, one obtains [14], [1], [2] $\mathcal{J}(\omega)={\omega^{2}}{\left(\pi\hbar\varepsilon_{0}c^{2}\right)^{-1}}\mathbf{p}\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{a},\mathbf{r}_{a}\right)\right]\cdot\mathbf{p}$ as a consequence of the fundamental integral relation 31 of [15]; here $\mathbf{r}_{a}$ is the position vector of the emitter and $\mathbf{p}$ is its transition dipole moment. If, instead, both reservoirs are initially in thermal states at the same temperature $T_{0}$, the reduced dynamics of the quantum emitter can still be recast in terms of a single zero-temperature bosonic effective reservoir with both positive and negative frequencies and a temperature-dependent equivalent spectral density equal to $\theta(\omega,T_{0})\mathcal{J}(|\omega|)$ where $\theta(\omega,T_{0})=\frac{1}{2}\text{sign}(\omega)\left[1+\coth\left(\frac{\beta_{0}\hbar\omega}{2}\right)\right]$ and $\beta_{0}=1/(k_{B}T_{0})$. Relations of this type are widely used in the Langevin- noise literature; however, the conditions under which they remain valid are not always clearly stated. Indeed, when the two reservoirs are prepared at different temperatures, the knowledge of the dyadic Green function of the dielectric objects at the position of the emitter is insufficient to determine the emitter dynamics. Nevertheless, the reduced dynamics of the emitter can still be described by a single bosonic reservoir initially in the vacuum state with an equivalent spectral density equal to the sum of the temperature-dependent medium-assisted and the temperature-dependent scattering-assisted spectral densities [2]. Scenarios in which two reservoirs are prepared in thermal states at different temperatures arise in several contexts [18] and may model different experimental settings, such as cavities driven by thermal photon baths in polaritonic chemistry (e.g., [19, 20]), and cryogenically cooled cavities (Fabry-Pérot, nanoplasmonic structures, photonic crystal waveguides, etc.) in the presence of thermal photons (e.g. [21, 22, 23]). In general, if the initial states of the two reservoirs are not thermal, it is no longer possible to represent their action through an equivalent single reservoir.
@@
Although the approach introduced in [1], [2] can handle an electromagnetic environment prepared in an arbitrary initial quantum state, it is limited to a single quantum emitter. In the present article, we lift this restriction and generalize this approach based on the modified Langevin noise formalism to a collection of quantum emitters with arbitrary orientations of their transition dipole moments. As we did for the single quantum emitter [1], we apply the emitter-centered mode technique to reduce the number of degrees of freedom of the electromagnetic environment (e.g. [24], [25]). In particular, we show that the reduced dynamic of the quantum emitters depends only on the values of the dyadic Green function at the quantum emitter positions whenever the initial state of the entire system is a product state and the initial states of the medium-assisted and the scattering-assisted reservoirs are thermal states with the same temperature. When these conditions are not satisfied, the reduced dynamics of the quantum emitter depends on the generalized spectral densities of the two individual reservoirs, which are quadratic functionals of the dyadic Green function of the dielectric objects.
@@
We proceed as follows. In Section II, we apply the modified Langevin noise formalism to model multiple quantum emitters interacting with a dispersive electromagnetic environment. In Section III, we use emitter-centered modes to reduce the number of degrees of freedom of the electromagnetic environment. In Section IV, we introduce the medium- and the scattering-assisted spectral density matrices. In Section V, we propose a surrogate bosonic environment to describe the reduced dynamics of quantum emitters when the initial quantum state of the entire system is a product state and the medium-assisted and scattering-assisted bosonic reservoirs are initially in thermal quantum states with different temperatures. In Section VI, we apply our model to a system of two quantum emitters interacting with a nanophotonic environment. We consider two scenarios, a metallic nanosphere and a metallic nanostructure composed of two rods and a disk. Specifically, we explore various aspects of entanglement relaxation, revivals, and generation. In Section VII, we give a summary and discuss our conclusions.
@@ II Model
A collection of $N$ quantum emitters mutually interacts and couples to finite-size dielectric objects of arbitrary shape embedded in an unbounded space. We denote by $\mathbf{r}_{j}$ the position vector of the $j$-th quantum emitter with $j=1,2,\ldots,N$. We assume that the dielectrics are linear, isotropic, and dispersive in time. We denote by $V$ the region occupied by the dielectric objects and by $\varepsilon_{\omega}(\mathbf{r})$ their relative dielectric permittivity in the frequency domain. The dispersive dielectrics, together with the electromagnetic field, constitute the electromagnetic environment of the quantum emitters.
@@ II.1 Hamiltonian
The Hamiltonian of the entire system, a collection of $N$ quantum emitters plus the electromagnetic environment, is
@@
where $\hat{H}_{A}$ is the bare Hamiltonian of the quantum emitters, assumed to be mutually isolated, $\hat{H}_{E}$ is the bare Hamiltonian of the electromagnetic environment and $\hat{H}_{I}$ describes their interaction.
In the multipolar coupling scheme (Power–Zienau–Woolley picture) and within the dipole approximation, the interaction Hamiltonian $\hat{H}_{I}$ reads
$\hat{H}_{I}=-\sum_{j=1}^{N}\hat{\mathbf{d}}_{j}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j}),$ (2)
where $\hat{\mathbf{E}}(\mathbf{r})$ is the electric field operator and $\hat{\mathbf{d}}_{j}$ is the electric dipole moment operator of the $j$-th emitter located at $\mathbf{r}_{j}$. We assume that the dipole moment operator of the $j$-th emitter couples through a fixed polarization direction $\mathbf{u}_{j}$ as in ref. [25], so that $\hat{\mathbf{d}}_{j}=\hat{\mu}_{j}\mathbf{u}_{j}$ where $\hat{\mu}_{j}$ is the corresponding transition dipole moment operator. More generally, one may retain up to three orthogonal polarization components per emitter, $\hat{\mathbf{d}}_{j}=\sum_{a}\hat{\mu}_{ja}\mathbf{a}$ where $a=x,y,z$, and $\hat{\mu}_{jx},\hat{\mu}_{jy},\hat{\mu}_{jz}$ are the corresponding transition-dipole moment operators.
II.2 Diagonal Form of $\hat{H}_{E}$
The bare Hamiltonian of the electromagnetic environment $\hat{H}_{E}$ accounts for the electromagnetic field, the polarization currents of the dielectric objects, and their interaction. The modified Langevin noise formalism provides a straightforward route to diagonalize it [15].
We express the electric field operator in the Schrödinger picture as $\hat{\mathbf{E}}(\mathbf{r})=\int^{\infty}_{0}d\omega[\hat{\mathbf{E}}_{\omega}(\mathbf{r})+h.c.]$ where $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ is its monochromatic component in the Heisenberg picture. Within the modified Langevin noise formalism, the electric field operator $\hat{\mathbf{E}}_{\omega}(\mathbf{r})$ has two contributions, the medium-assisted contribution $\hat{\mathbf{E}}^{(M)}_{\omega}(\mathbf{r})$ and the scattering-assisted contribution $\hat{\mathbf{E}}^{(S)}_{\omega}(\mathbf{r})$,
@@ II.2 Diagonal Form of $\hat{H}_{E}$
Let $\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})$ denote the solution of a homogeneous equation
$\left(\nabla_{\mathbf{r}}\times\nabla_{\mathbf{r}}\times-k_{\omega}^{2}\varepsilon_{\omega}\right)\mathbf{F}_{\omega\mathbf{n}\nu}=0,$ (7)
satisfying the asymptotic boundary condition
$\mathbf{F}_{\omega\mathbf{n}\nu}(\mathbf{r})\underset{r\rightarrow\infty}{\approx}e^{ik_{\omega}\mathbf{r}\cdot\mathbf{n}}\mathbf{e}_{\mathbf{n}\nu},$ (8)
@@
where $\hat{g}_{\omega\mathbf{n}\nu}$ is the monochromatic bosonic operator that describes the fluctuation of the radiation incoming from infinity and scattered by the dielectric object. Here $o_{\mathbf{n}}=(\theta_{\mathbf{n}},\phi_{\mathbf{n}})$ are the polar angles of the unit vector $\mathbf{n}$}, $do_{\mathbf{n}}=\sin\theta_{\mathbf{n}}d\theta_{\mathbf{n}}d\phi_{\mathbf{n}}$ is the differential of the solid angle and the integration is over the full solid angle, $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$.
@@
${\left[\hat{\mathbf{f}}_{\omega}(\mathbf{r}),\hat{\mathbf{f}}_{\omega^{\prime}}^{\dagger}\left(\mathbf{r}^{\prime}\right)\right]}$ $=\delta\left(\omega-\omega^{\prime}\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)\mathcal{I},$ (11a)
@@
where $\delta\left(o_{\mathbf{n}}-o_{\mathbf{n}^{\prime}}\right)=\delta\left(\theta_{\mathbf{n}}-\theta_{\mathbf{n}}^{\prime}\right)\delta\left(\phi_{\mathbf{n}}-\phi_{\mathbf{n}}^{\prime}\right)/\sin\theta_{\mathbf{n}}$. Together with the expressions of the medium-assisted electric field $\hat{\mathbf{E}}^{(M)}_{\omega}$ and of the scattering-assisted electric field $\hat{\mathbf{E}}^{(S)}_{\omega}$, these commutation relations ensure the canonical commutation relations for the electromagnetic field and the bath oscillator fields that describe dielectric objects [15].
@@
The fundamental integral identity [14, 15]
@@
This identity was first validated numerically by D. Y. Na and coworkers [14].